my father is 23 years older than me In 15 years ge will be triple my age today. How Old am I

I keep getting stuck in the equation I make the father X and the child Y
therefore x=y+23, after 15 years x = 3 x y
then I do this but I get in a mess
(x+15) = 3 x (y +15)
then I get completely lost, tried to put it in table form, no answer is the same.
Please help, there is nothing online that relates to this type of question and I know its simple but I am 15 and struggle with math questions

Found 2 solutions by jorel1380, Alan3354:
Answer by jorel1380(3719)

Let x be your age today. Then your father's age is x+23. So:
3(x)=x+23+15
3x=x+38
2x=38
x=19
x+23=42
You are 19; your father is 42. ☺☺☺☺

Answer by Alan3354(69443)

my father is 23 years older than me In 15 years ge will be triple my age today. How Old am I
----------
In 15 yrs:
F = 3*(Y-15)
F = Y + 23
---
3Y-45 = Y + 23
2Y = 68
YOU will 34 in 15 years
-------
Now you're 19, he's 42
He'll be 57 in 15 years, = 3*19